of performance of an ... An ideal refrigerator does 175.0 J of work to remove 780.0 J of heat from its cold compartment during each cycle. What is the refrigerator's coefficient of performance? Tries 0/20

How much heat per cycle is exhausted to the kitchen?
Answer #1

here,

the work done,W = 175 J

the heat removed from the cold chamber during each cycle , Qc = 780 J

the COP = W /Qc

COP= 175 /780

COP = 0.224

the heat exhausted to the kitchen , Qh = W + Qc

Qh = 175 J + 780 J

Qh = 955 J